### v2

# EMMOE: A Comprehensive Benchmark for Embodied Mobile Manipulation in Open Environments

Abstract
Developing autonomous home robots controlled by natural language has long been a pursuit of humanity. While advancements in large language models (LLMs) and embodied intelligence make this goal closer, several challenges persist: the lack of a unified benchmark for more complex robot tasks, limited evaluation methods and metrics, data incompatibility between LLMs and mobile manipulation trajectories. To address these issues, we propose Embodied Mobile Manipulation in Open Environments (EMMOE), a benchmark that requires agents to interpret user instructions and execute long-horizon everyday tasks in continuous space. EMMOE seamlessly integrates high-level and low-level embodied tasks into a unified framework, along with three new metrics for more diverse assessment. Additionally, we collect EMMOE-100, which features in various task attributes, detailed process annotations, re-plans after failures, and two sub-datasets for LLM training. Furthermore, we design HomieBot, a sophisticated agent system consists of LLM with Direct Preference Optimization (DPO), light weighted navigation and manipulation models, and multiple error detection mechanisms. Finally, we demonstrate HomieBot’s performance and evaluations of different models and policies. All datasets, models, and codes can be found in our Project Website.
1 Introduction
Figure 1: Data example in EMMOE-100. A key feature of EMMOE-100 is the emphasis on the reasoning process and interleaved execution. In the shown task, the agent must check the fridge first. Otherwise, even if the agent finally gets a banana in the kitchen, it will not be considered as a success.
Developing autonomous robots capable of performing various daily tasks through a single instruction has been a long-standing goal. To achieve this goal, robots need to understand natural language instructions, make feasible plans, perceive and interact with dynamic environments, and equip with powerful navigation and manipulation skills. Typical methods like imitation learning (IL) [1] and reinforcement learning (RL) [2] primarily focus on task-specific policies, but are always limited to short-horizon tasks and struggle to generalize to new tasks. Task and Motion Planning (TAMP) treats long-horizon mobile manipulation tasks as hybrid discrete-continuous search problems [3] and addresses with a hierarchical architecture [4]: High-level task planning in discrete task space, low-level motion planning in continuous action space, and interleaved execution between two layers. However, the scope of TAMP remains limited. Despite various extensions [5, 6, 7], it still requires specific goal states and detailed scene configurations. The complexity and dynamism of real-world environments, and vague user descriptions make it highly challenging to meet these requirements.
In recent years, the rapid development of LLM [8, 9] and embodied intelligence [10, 11] has made this pursuit possible. The scope of each layer in TAMP has been largely broadened and spawns various embodied tasks driven by language and vision. In high-level embodied tasks [12, 13], LLMs have shown exceptional performance and powerful generalization capabilities. Advanced prompting techniques like Chain-of-Thought (COT) [14] have further enhanced the logical reasoning abilities of LLMs. Visual Language Models (VLMs) [15] enable agents to process visual inputs and understand current environments. Large Multi-modal Models (LMMs) [16] extend the application of embodied agents to real-world scenarios. The most recent world models [17] and spatial models [18] allow agents to more accurately perceive scene information and spatial relationships. In low-level embodied tasks, the emphasis of models has gradually shifted from single skill with specific objects [19] to single skill with open-vocabulary objects [20], then further to general models [21], such as Visual Language Navigation (VLN) [22] and Visual Language Action (VLA) [23] models.
However, several problems remain unresolved: 1) Lack of a comprehensive task and benchmark. Although significant progress has been made in various embodied tasks, there is still a gap between the current tasks and the envisioned language-driven intelligent robots. Meanwhile, existing embodied tasks always operate in isolation, neglecting the mutual influence caused by interleaved task execution. By integrating different high-level and low-level embodied tasks, robots can achieve more advanced capabilities while enabling a unified evaluation of various embodied tasks. Each layer will constrain and influence the others, working collaboratively to accomplish the final task. 2) Inadequate evaluation methods and metrics. Embodied task planning involves causal dependencies between each step, and solutions are not absolute, thus making evaluations based solely on individual steps or the final state insufficient. Additionally, current evaluation methods rely heavily on simulators or PDDL files, which also limits the real-world deployment and application of agents. Furthermore, how to make more fine-grained evaluations of the entire agent system remains a challenge. 3) LLM grounding problems. Although LLMs excel in commonsense reasoning, they need to be grounded in current environments to produce realistic and practical outputs. Furthermore, due to the uncertainties and dynamic changes in the real world, LLMs must be able to make timely adjustments based on real-time feedback. However, the incompatibility between the conversation data required for LLMs and the trajectory data required for robotics increases the difficulty of grounding.
To advance the development of intelligent autonomous robots, we propose EMMOE as an open challenge, which requires agents to interpret user instructions and execute long-horizon everyday tasks in continuous space. Besides, we manually collect EMMOE-100, the first daily task dataset featuring various task attributes, detailed process annotations, analyses of each output, re-plans after failures. We also build Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO) [24] sub-datasets to facilitate the alignment of LMM capabilities with specific embodied tasks. Finally, we introduce HomieBot, a sophisticated agent system that integrates both high-level and low-level models, as well as multiple error detection and adaptation mechanisms. An example of EMMOE challenge and EMMOE-100 tasks is shown in Fig.1.
Table 1: Dataset Comparisons. EMMOE-100 is the first dataset to integrate mobile manipulation tasks with embodied task planning, decomposing long mobile manipulation trajectories into discrete actions then executed by low-level policies in continuous space.
Benchmark Low-level Policy Selection Task Planning Manipulation Navigation Procedure Annotations Re-plan LMM Trainable Format COT Analysis Open-ended Instructions DPO Sub-dataset
OVMM Continuous Continuous
BEHAVIOR-1K Continuous Continuous
ALFRED Discrete Discrete
Octopus Discrete Discrete
Habitat-LAB 2.0 Continuous Continuous
VirtualHome Discrete
ManiSkill-2 Continuous Continuous
Grutopia Continuous Continuous
EMMOE-100 Continuous Continuous
In particular, our paper makes the following contributions:
  • We propose EMMOE, the first unified benchmark for both high-level and low-level embodied tasks with three novel metrics for more advanced evaluation.
  • We collect EMMOE-100, the first everyday task dataset featuring COT outputs, diverse task designs, re-plan processes, with SFT and DPO sub-datasets.
  • We design HomieBot, a sophisticated agent system which integrates models at different levels, multiple error detection and adaptation mechanisms.
2 EMMOE Benchmark
2.1 Problem Statement
EMMOE requires that robots explore environments and perform various open-vocabulary mobile manipulation tasks based solely on language instructions and sensor observations. More specifically, it combines embodied task planning, embodied decision making, visual language navigation in continuous environments, and language-conditioned manipulation, which requires highly on both level of models and the design of agent systems.
2.2 EMMOE-100 Dataset
By controlling Fetch Robots [25] in Habitat-Lab 2.0 [26], we collect EMMOE-100, a dataset consists of 100 complex everyday tasks. We sample 100 different scenarios from Replica Challenge [26] to build simulation environments. In each scene, we’ll first design a daily mobile manipulation task, then manually control a Fetch robot to complete the task in continuous space and decompose execution trajectories into discrete subtasks. Each subtask consists of an executable action, a target, and a low-level model selection, finally we obtain 966 subtasks in total. We also annotate each subtask with four first-person view images and detailed reasoning processes. Moreover, we intentionally design some failed subtasks and provide re-plans to enhance dataset robustness. To alleviate grounding problems, we construct SFT and DPO sub-datasets, which will be introduced in Section 4.1.
To enhance task diversity and better align with human demands, we design tasks with five different attributes: Short-horizon tasks like pick something and place it somewhere. Long-horizon tasks which consist of at least ten subtasks. Open-ended tasks that allow multiple results and solutions. Logical tasks that provide vague descriptions and require logical reasoning. Human-style tasks are described in a natural conversation style. One task can possess multiple attributes as some of these attributes are not contradictory. Table 1 shows detailed comparisons with other mobile manipulation and embodied task datasets. We also provide detailed task statistics in Appendix B.
2.3 Evaluation Metrics
The most fundamental metrics in embodied task planning are Success Rate (SR) and Goal-Condition Success (GC) [27]. SR measures the proportion of successful trajectories, while GC is the ratio of goal conditions achieved at the end of a trajectory. A trajectory is considered successful only if GC reaches 100%. However, GC focuses only on the final result and relies on pre-defined state goals, thus failing to meet the requirements of our EMMOE tasks, which require fine-grained and language-based evaluations. Although some studies [13] conduct more fine-grained evaluations, they overlook the flexibility and coherence in embodied task planning and still rely on abstract terms. The success of an individual step may not contribute to the final success, and an output that differs from the ground truth but can complete the task in an alternative way should not be considered incorrect. Furthermore, fine-grained evaluation of the entire agent system remains a challenge. To overcome these limitations and provide more diverse evaluations, we propose the following new metrics. All details about definitions and visible calculation examples can be found in Appendix C.
Task Progress
To better measure the task execution process and the interrelations among subtasks, we propose Task Progress (TP), which is calculated as follows:
$TP=\max_{k_{i}\in K_{T}}\left(\frac{\text{len}(k_{i}^{\text{check}})}{\text{len}(k_{i})}\right)$ (1)
A keypath is defined as an ordered node set of all necessary subtasks required to complete a task, $k_{i}$ is the $i$-th keypath in the keypath set $K_{T}$ for task $T$, each task is assigned with several keypaths, representing different ways to complete the task. We strictly match the execution trajectory with the subtask nodes in $k_{i}$ in sequential order. Once the node in $k_{i}$ is successfully matched, it will be added to another ordered set $k_{i}^{\text{check}}$, then the ratio between the length of $k_{i}^{\text{check}}$ and the length of $k_{i}$ will be recorded. This process will be repeated for all keypaths in $K_{T}$, and the highest ratio will become the TP value of the trajectory. Only if TP reaches 100%, the trajectory will be considered successful. TP considers both the flexibility of the execution process and the relationships between every step. The way of using natural language and execution results to evaluate also simplifies new task design and enables evaluation in real-world scenarios, where writing PDDL files is impractical.
Success End Rate
A fully autonomous robot should be able to actively terminate the execution at a proper moment. Otherwise, even if the task is already done, the robot may continue running and get stuck in an endless loop. Therefore, we propose Success End Rate (SER) to evaluate whether the agent has the ability to understand its current situation and reasonably determine the appropriate timing for task termination, the calculation method is as follows:
$SER=\frac{\text{len}(S)}{\sum_{t\in M}\text{count}_{t}(\text{end})}$ (2)
$t$ represents a single trajectory and $M$ is the set of trajectories for all tasks, $\text{count}_{t}(\text{end})$ equals 1 if $End$ is the final action of $t$ or 0 otherwise. $S$ is the set of successful trajectories, of which TP equals 100%. Then SER is calculated as the ratio of the number of successful trajectories to the number of trajectories that the agent deemed successful. Once SER reaches a certain threshold or even 100%, auxiliary methods or metrics are no longer needed to calculate SR.
Success Re-plan Rate
Execution failures are common cases in the real world, especially in unfamiliar environments, which makes the ability to quickly adjust from failures and continuously adapt to new environments a crucial skill. To measure the adaptation and generalization abilities of the agent, we propose Success Re-plan Rate (SRR), which is calculated as follows:
$SRR=\frac{\sum_{t\in S}\text{count}_{t}(\text{replan})}{\sum_{t\in M}\text{count}_{t}(\text{replan})}$ (3)
$\text{count}_{t}(\text{replan})$ is the number of re-plans in trajectory $t$, other symbol definitions are the same as SER. SRR represents the effectiveness of re-planning and adaptability of the agent. When SRR reaches 100%, it indicates that the agent can adapt to all failures and then successfully complete the task.
3 HomieBot
3.1 Overview
In this section, we will introduce how HomieBot accomplishes EMMOE tasks. HomieBot employs a hierarchical framework with communication mechanisms for interleaved execution. High-Level Planning (HLP) deals with embodied decision making and planning adaptation, while Low-Level Execution (LLE) translates subtasks into continuous low-level controls and provides feedback to HLP. We will describe HLP in Section 3.2 and LLE in Section 3.3. A system overview is shown in Fig.2.
Figure 2: Overview of HomieBot. HomieBot leverages a hierarchical framework to handle long-horizon tasks: High-Level Planning decomposes tasks into manageable actions, Low-Level Execution accomplishes received actions and provides real-time feedback.
3.2 High Level Planning (HLP)
A long trajectory will be decomposed into several subtasks, the agent must continuously interact with the environment and adjust plans based on real-time feedback to ensure generated subtasks are practical. We design elaborate input and output instructions to facilitate dynamic adjustments during execution. Video-LLaVA [28] is selected as our planner model $M$ and fine-tuned with SFT and DPO sub-datasets, which will be described in Section 4.1.
Multi-modal Instruction
To help the LMM better understand current situations, the format of the input instruction $I$ is as follows:
$I=\{o_{1\sim 4},s,T,inv,h,f\}$ (4)
In the visual component, four first-person view images $o_{1\sim 4}$ correspond to four directions respectively. In the textual component, system information $s$ and user task $T$ remain constant throughout the conversation, reminding the agent of its responsibility. Feedback $f$ indicates the status of the last execution and detailed error information if failed, it will also be used to update other parts in $I$. Inventory $inv$ reflects the items currently held by the agent, primarily to prevent the generation of illogical actions, it is updated based on both $f$ and the type of the last action. Execution history $h$ logs all previous subtasks and their results. Once receiving $f$, the last subtask and its result will be logged in $h$. Besides, to better align with real-world scenarios, we prohibit directly inputting background information into the LMM (e.g. raw object data, Bird’s Eye View images etc.). The planner must explore the environment and enhance its intrinsic capabilities to generate more reasonable outputs.
Json-format Output
Considering that different low-level policies may require different information formats and to facilitate the replacement and maintenance of each module, we define our output in the following uniform format:
$O=M(I)=\{A,S,m\},S=\{\texttt{action},\texttt{target}\}$ (5)
$A$ represents the analysis of each output, which is inspired by works like CoT [14]. Before generating final outputs, planner model $M$ is expected to summarize previous executions and current situations, analyze what to do next, and propose the subsequent subtask $S$. To ensure the feasibility of the output, action can only be chosen from the available action list. Similarly, $m$ which represents the selected low-level models or policies, is also restricted to a given model list. target can be either an object or a spot, it should be observable in the provided images and deemed necessary to complete the task.
3.3 Low Level Execution (LLE)
LLE will convert $S$, $m$ and $inv$ from HLP into precise model-calling instructions. Error detection will be applied at different stages to monitor the whole process. Once the execution is completed or failed, environmental images and feedback will be sent back to HLP. We set up six skills based on the support of the simulator(see Table E1). Since the required information varies from models and would significantly impact the model performance, we establish two distinct settings to ensure fairness.
Execution With Background Information
More specifically, execution with background information means that the selected model needs precise position and state information of the target. As M3 [29] shows exceptional performance in all skills when utilizing background information in Habitat, we choose it as the unique model choice $m$ in this setting. To make M3 adapt to our task requirements, we implement a name mapping for $target$ and adjust its original setting to better align with requirements of our tasks. In addition to text and image data, LLE also captures the execution process of each step and the entire trajectory data in video format. This means that HomieBot has the potential to bridge the gap between robot data and LMM data as the entire execution process is fully automated and annotated, users only need to set up the scene and input instructions. The video data can be utilized for IL in robotics, while the text and image data can be utilized for LMM training.
Execution Without Background Information
Without background information means that the agent can only rely on the information captured by its sensors and the intrinsic abilities of low-level models to complete the task. As shown in Table E2, we set two manipulation models and two navigation models. For manipulation, RT-1-X [30] is used for $Pick$ and $Place$, while Octo [31] is set for $Open$ and $Close$. For navigation, NoMaD [32] specializes in image navigation and is suitable when $target$ is a spot or large object. PixNav [33] excels in pixel-level and object navigation, making it ideal when $target$ is a detectable object. As the deployment of robots in the real world always demands high real-time performance and is constrained by hardware limitations, we prefer to choose lightweight models rather than the currently popular VLA models to prevent the system from becoming too burdensome. Additionally, breaking down long-horizon tasks into action primitives would also reduce the performance requirements of low-level models. Compared to general-purpose end-to-end models, specialized lightweight models can complete the action while reducing time costs.
Error Detection
To facilitate communication with HLP and provide more detailed error information, we further classify common errors into four main types and several subtypes. Logical error L1: The agent’s hands are already full but still attempts to pick/open/close; L2: The agent holds nothing but attempts to put; L3: The agent attempts to pick/put the object in a closed container; L4: The agent attempts to interact with a non-interactive object. Distance error D1: The agent stands too far and is unable to reach the target; D2: The agent is too close to the target and its arm is hindered from properly extending during interaction. Format Error F1: The output action or model is not in the available list; F2: The output target does not exist in the current scene or can not be recognized by low-level models. Execution Error E1: The limited capabilities of the low-level models or policies cause the failure; E2: Failed execution may result in the inventory information being accidentally updated. Furthermore, we conduct multiple phases of error detection during the whole process to guarantee the executions. More classification and detection details are given in Appendix E.
4 Experiments
4.1 Data Augmentation
SFT Augmentation
Previous work[34] has shown that a standardized data format would significantly enhance model training and evaluation. Therefore, we write a uniform script to convert the original EMMOE-100 data into fixed-format conversation data. During this process, all failed subtasks will be skipped as they are treated as junk data for the SFT dataset, and we initially obtained 930 SFT data in this way, which is still insufficient for LLM training. To expand the dataset, we use GPT-4o [35] to regenerate text descriptions of tasks and the analysis of each subtask for three times. This approach not only enhances the diversity of instructions, allowing the LLM to adapt to different user input styles, but also helps to avoid introducing additional inaccuracy or inconsistency. Finally, we obtain 3,720 SFT data in total. The relevant code and data samples are available in Appendix F.1.
DPO Augmentation
DPO [24] training has a strict requirement for data format, which must include prompt, chosen and rejected. For the $i$-th subtask and its input instruction $I_{i}$, if the execution of output $O_{i}$ fails but the next output $O_{i+1}$ succeeds after re-plan, we will choose $I_{i}$ as the prompt, $O_{i}$ as the rejected and $O_{i+1}$ as the chosen. Although this approach aligns well with the concept of preference data, the proportion of re-planned data is relatively low. Thus, we utilize following methods to construct new DPO data. Order Change: We shuffle the order of successful subtasks, treating successful output $O_{i}$ as chosen and $O_{i+1}$ as rejected. This approach aims to help LLMs learn the logical relationships between subtasks, particularly the optimal sequence of actions. Action Change: To standardize the planner model’s output and reduce responses outside the action list, we replace actions in subtasks with non-standard names or actions outside the available list. Model Change: To enable the LLM to own the ability to select the appropriate low-level model for a given scenario, we replace the model choice with models of the same type in the model list. As a result, we get 10,104 DPO data in total. More processing flows and data samples are provided in Appendix F.2.
4.2 Model Training
We select 90 tasks from EMMOE-100 as our training tasks. Using the methods described in Section 4.1, we obtain 3,316 SFT training data and 8,984 DPO training data in total. Then we select Video-LLaVA-7B [28] as our base model and conduct a two-stage training process. In the first stage, we fine-tune the base model with a learning rate of 5e-4 on 4$\times$×NVIDIA A40. In the second stage, we align the fine-tuned model with DPO and train with a learning rate of 5e-6. To prevent catastrophic forgetting and maintain the intrinsic model capability, LoRA [36] is applied in both stages, with LoRA rank set to 128 and $\alpha$ to 256 in stage one, and LoRA rank set to 8 and $\alpha$ to 8 in stage two.
4.3 Setup
Metrics
In addition to SR, TP, SER and SRR introduced in Section 2.3, we also choose Path Length Weighted SR (PLWSR)[27] as one of our evaluation metrics. PLWSR is defined as SR×$\times$(length of successful trajectory) / $max$(length of expert trajectory, length of successful trajectory) and measures the ability gap between the agent and the expert in successful trajectories.
Baselines
High Level Planner: In baseline planner selections, GPT-4o [35] and Gemini-1.5-Pro [37] are the most popular and common closed-source models. The reasoning model o1 [38] is famous for its powerful reasoning abilities and can effectively handle complex inference tasks. As for open-source models, we choose Qwen2-VL-7B [39] and MiniCPM-V 2.6 [40], which perform well in various multi-modal tasks and have similar model sizes to HomieBot. GPT-4o, Gemini-1.5-Pro and o1 can be easily integrated into HomieBot after minor adjustments to format requirements. By leveraging the in-context learning abilities and providing output examples for each inference, the other two models can also be deployed in our system. Low Level Executor: We extract individual skills from M3 [29] and modify their implementations. Original skills require the initial and final states of the object. We map the object name to obtain specific background information and select the nearest object. Additionally, robotic arms will be reset after each execution to enhance the success rate. We also pass all environmental state information between executions to ensure environmental consistency. We use single NVIDIA A40 to run both models and provide more details in Appendix H.1.
Evaluation Benchmarks
All tasks in EMMOE-100 will be used for evaluation, and the remaining ten untrained tasks will serve as our test set. Each task is executed three times with a maximum step limit of 20 each time, the average execution results will be used for the final calculation.
4.4 Results
We begin with a general evaluation since all data are unseen to baseline models. As shown in Table 2, the DPO version of HomieBot achieves the best performance in SR, PLWSR, TP and SER metrics. o1 also demonstrates excellent performance in our tasks and surpasses the SFT version in some metrics. Additionally, it is evident that for open-source models of similar size, even state-of-the-art LMMs like Qwen2-VL-7B [39] and MiniCPM-V 2.6 [40] struggle in EMMOE tasks without additional training.
For SER, though the DPO version still performs best, the improvement is not so obvious. This phenomenon should be attributed to the nature of SER, which reflects the model’s ability to correctly determine when a task is completed and should be terminated. It is less influenced by format requirements and low-level executions, but relies more on the model’s inherent reasoning ability. The strong reasoning capabilities of GPT-4o [35], Gemini-1.5-Pro [37] and o1 [38] enable them to effectively decide when to end a trajectory. Moreover, as we observed during experiments, o1 is more likely to “give up” compared with other models. After several failed attempts, o1 would judge that the current task is infeasible and directly terminate task execution, while other models would continue to explore. This tendency results in relatively lower SER, and also affects TP to some extent.
For SRR, o1 performs best and SFT version performs better than DPO version. Since SRR reflects the model’s ability to adapt to environments and adjust from failure, this result suggests that o1 can better leverage feedback information to make more effective re-plans. Besides, it could also be relevant with the limitations of the DPO method [41]. Although DPO brings unparalleled advantages in training efficiency and convenience, it compromises the model’s generalization and transferability to certain extent. Therefore, we further evaluate HomieBot separately in training and test set. As we can observe in Table 3, while DPO version performs best on all metrics in the training split, it only outperforms SFT version on SER in the test split. Additionally, DPO version shows a significant decline on SRR. This observation further confirms that the DPO method introduces certain generalization issues.
Notably, SER remains stable for both versions across the training and test splits, which further demonstrates that SER is more related to the model’s inherent judgment ability, and our specialized handling of $End$ during dataset construction has enhanced this ability (See in Appendix F.2).
Table 2: Performance comparison of different models on EMMOE-100 tasks. The highest values for each metric are highlighted in bold.
Model SR PLWSR TP SRR SER
Qwen2-VL-7B[39] 1.00 0.50 16.55 0.59 25.00
MiniCPM-V 2.6[40] 0.67 0.57 14.45 0.06 40.00
GPT-4o[35] 13.33 10.51 29.79 3.57 49.38
Gemini-1.5-Pro[37] 17.33 14.79 38.03 3.39 55.91
o1[38] 28.67 24.11 44.52 13.80 38.57
HomieBot-7B (SFT) 27.67 20.88 50.27 9.23 53.90
HomieBot-7B (SFT+DPO) 30.30 24.66 51.39 8.72 60.81
Table 3: Performance comparison of HomieBot on the training and test split. The highest values for each metric are highlighted in bold.
Model Train split Test split
SR PLWSR TP SRR SER SR PLWSR TP SRR SER
HomieBot (SFT) 28.52 21.49 50.16 9.59 53.85 20.00 15.36 51.19 6.55 54.55
HomieBot (SFT+DPO) 31.84 25.82 52.29 9.69 60.71 16.67 14.36 43.39 3.08 62.50
4.5 Analysis
To further explore the reasons for the overall low success rate and demonstrate how HomieBot can be used to simultaneously evaluate both HLP and LLE, we conduct a detailed analysis based on the results in Section 4.4. Using the error classification in Section 3.3 and the recorded feedback, we collect all errors that occurred during experiments. To identify which errors are acceptable and solvable and which are the primary causes of failure, we further classify the collected errors according to whether they appear in successful or failed trajectories, the results are shown in Figure 3.
Error Analysis
Except for $E1$ and $E2$ error that come from low-level models, each error type corresponds to different capabilities of LMMs. In failed trajectories, the predominant error type across all baseline models is $F2$ error. This suggests that the primary obstructive factors are physical grounding failures and model hallucinations. In practical execution processes, we observe that even models are already told the object doesn’t exist or can’t be recognized, they may still produce inappropriate outputs or repeat mistakes after several steps. This issue has been significantly improved in our models, which also highlights the significance of LMM-trainable format data. With a small amount of augmented data, LMM can build up a general understanding of current environments, enabling outputs to be grounded and compatible with low-level models.
Besides, the proportion of failed executions for two open-source models is relatively low, indicating that most subtasks are successfully completed, which seems to conflict with the very poor SR. Based on our observations, since EMMOE includes numerous complex and long-horizon tasks, execution histories often become lengthy. When the model’s understanding ability is insufficient, it may fail to fully understand or even forget previous execution contents, ultimately resulting in meaningless outputs. Although these subtasks can be successfully executed, they contribute nothing to the final task, and even worse, they will consume remaining steps and fasten task termination. In successful trajectories, the most common error is $D1$ error. This indicates that even when the model’s spatial perception ability is insufficient, it can be adjusted through feedback information. Typically, after a $D1$ error occurs, the model will output $Go\ to$ action based on the feedback, effectively resolving this error. We conduct more detailed case studies in Appendix I.
Figure 3: Error Statistics. The left and right figures depict the proportion of each error type of each model in successful and failed trajectories respectively. Additionally, we indicate the proportion of total execution failures next to each model’s name. Due to too few successful trajectories for Qwen2-VL and MiniCPM-V 2.6, their results will not be shown in the left figure. The full statistical data in digital counts are available in Appendix H.2.
Table 4: Results of LLE evaluations. P represents the proportion of single action errors. SR here represents an average value as each skill is attempted up to three times per execution.
Metrics Go to Pick Place Open Close
P 38.49 49.77 7.30 3.32 1.11
SR 45.32 22.45 40.97 43.13 36.45
LLE Evaluation
Comprehensive error types allow us to evaluate HLP and LLE separately. We further classify $E1$ and $E2$ errors based on action types and count total occurrences of each action, the calculation results are shown in Table 4. It is evident that $Pick$ action has a significantly lower success rate and the highest proportion of execution errors compared to other actions.
5 Conclusions and Discussions
We propose EMMOE, the first unified benchmark designed to evaluate both high-level planners and low-level policies. Then we present the collection and features of EMMOE-100 and propose three novel metrics to complement existing evaluation methods. Next, we introduce our HomieBot and illustrate how its two main components HLP and LLE function. In experiments, we demonstrate how to construct LMM-trainable SFT and DPO datasets and evaluate different levels of models. Finally, we conduct an in-depth analysis based on the detailed error information.
Limitations and Future Work
Limited actions and available space in Habitat restrict the scope of task design. Besides, standardized output will sacrifice certain information precision. The growing number of model inferences will also lead to additional time costs. Moreover, we choose to conduct evaluations solely in simulation, where all researchers are required to make assessments under the same conditions, thus ensuring optimal fairness and consistency. In the future, we’ll collect more tasks, design a more efficient system, and explore real-world evaluations.
Appendix
The appendix is structured as follows:
  • Related Work in Section A.
  • Dataset in Section B.
  • Metric Calculation in Section C.
  • High Level Planning in Section D.
  • Low Level Execution in Section E.
  • Data Augmentation in Section F.
  • Training Details in Section G.
  • Experimental Details in Section H.
  • Case Study in Section I.
Appendix A Related Work
A.1 Embodied Tasks and Benchmarks
As embodied agents and LLMs develop rapidly, many embodied tasks and benchmarks have emerged. In Embodied Question Answering (EQA) tasks, EQA-v1 [42], VirtualHome [43], MT-EQA [44], MP3D-EQA [45], IQUAD V1 [46], OpenEQA [47], HM-EQA [48], S-EQA [49] contains a variety of task range to evaluate logical reasoning abilities of LLMs. BLINK [50] for visual perception abilities of LMMs. In Vision-and-Language Navigation (VLN) tasks, R2R [51], R4R [52] and VLN-CE [53], SOON [54] evaluate LLM’s capabilities under different navigation settings. ALFRED [27] Behavior series [55, 56] focus on interactive household tasks OVMM [57] involves picking and placing any object in unseen environments. VLA-3D [58] for 3D semantic scene understanding and navigation. Common manipulation datasets include MT-Opt [59], VIMA [60], ManiSkill2 [61], Calvin [62], BridgeData-v2 [63], RH20T [64], Open-X [65], AgiBot World [66]. In mobile manipulation, RT series [67, 68] and Mobile ALOHA [69] exhibit strong capabilities. GRUTOPIA [70] takes human participation into account. Additionally, some benchmarks focus on cross-embodiments, like RoboMIND [71]. EmbodiedBench [72] try to evaluate LMMs together in high-levels and low-levels. Despite numerous benchmarks, a unified benchmark and relevant task is still missing. Traditional mobile manipulation uses IL to learn entire trajectories, complicating the evaluation of intermediate processes. In our work, we propose EMMOE, a holistic benchmark designed to assess both final results and the execution process.
A.2 LLMs For Robotics
LLM-driven embodied agents represent cutting-edge advancements in robotics. SayCan [73], Palm-E [11], LLM-Planner [74] and EmbodiedGPT [75] combine LLMs with complex embodied tasks. TAPA [12] and SayPlan [76] use visual modules for multi-room settings. Voyager [77], STEVE  [78], Smallville [79] and Octopus [80] use LLMs to choose pre-defined functions. L3MVN [81], ESC [82], SayNav [83] and VLFM[84] build frontier or semantic maps to navigate. ViNT [85] and NoMaD [32] focus on image navigation, PixNav [33] uses LLM to select target image pixel. GOAT [86] is a comprehensive navigation system. Navid [22] and Uni-Navid [87] focus on end-to-end navigation models. RT-2 [68] is the first Visual Language Action (VLA) model. RoboFlamingo [88] and OpenVLA [89] are open-source VLA models. Leo [90] focuses on multiple QA problems. Manipulate Anything [91] and Octo [31] are light models for arm control. ALOHA [92] improves action prediction through action chunking. RoboAgent [93] enhances object detection and generalization, and LCB [94] uses LLMs to generate implicit strategy goals. ManipLLM [95], VoxPoser [96], Rekep [18] combine environmental perception and task execution.
A.3 LLMs for Task Planning
Typical usages of LLM for task planing include treating LLM as a translator or a planner. There are also some studies combining LLMs with traditional PDDL [97, 98, 99, 100], in which LLM will be treated as a translator between the real-world and specific domain planner. But this method is limited by the performance of the domain planner and can’t leverage the powerful commonsense reasoning capabilities of LLMs to assist in planning. When LLM is treated as a planner, discrepancies between LLM’s outputs and real-world conditions always lead to execution failures. LLM-Planner [74] make a straightforward re-plan. Self-Refine [101] use single LLM as generator and evaluator. Reflexion [102] treats LLM as the Actor and the evaluator as the Critic. ViLA [103] utilizes GPT-4V [104] to obtain visual feedback. However, self-improvement relies heavily on prompt design and intrinsic capabilities of LLMs. If errors unrelated to planning occur, LLMs may struggle to self-correct. Inner Monologue [105] and RoCo [106] utilizes external collision detection and feedback. DoReMi [107] sets pre-defined constrains. Nevertheless, LLMs may make same mistakes in similar situations as the model weights are not changed. SayCan [73] trains a value function to consider both generated actions and their values. Remember [108] builds a memory module and retrieves similar state-action pairs. Retroformer [109] learns a retrospective model via policy gradient optimization. While RL-based adaptation mechanisms are able to adjust actions before execution, defining and training an effective value function or reward model is highly challenging. The recently popular DPO [24] algorithm greatly simplifies this process by requiring only a preference dataset. In our  HomieBot, we use DPO for model alignment, CoT [14] and self-reflection for decision-making. Additionally, error detection and feedback mechanisms are applied during low-level execution.
Appendix B Dataset
B.1 Data Collection
We first randomly sample episode information provided by Replica Challenge [26] to build the task scenario, then we use the Fetch robot to collect EMMOE-100 in Habitat-lab v0.2.3. To facilitate data collection, we modify the original interaction script, and new interface can be seen in Fig. B1. The interface provides both third-person and first-person view observation to facilitate data collection, third-person observation is used to assist with collection, only first-person observation with 256*256 resolution will be saved. Notably, we only use the scene information to collect environment data, other functions and metrics in Replica Challenge are irrelevant to our work.
Figure B1: Data collection interface in Habitat-lab v0.2.3. Third-person observation in the left is used to facilitate data collection, only first-person observation with 256*256 resolution in the right will be saved.
We also show the modified code clip, once a single subtask is finished, we can conveniently save relevant information by pressing the keyboard.
def save_first_view_images():
directions = [’left’, ’back’, ’right’, ’front’]
global h_cnt
h_cnt += 1
for i in range(4):
for j in range(19):
base_action = [0, 1]
name = base_action_name
args = {base_key: base_action}
result = step_env(env, name, args)
use_ob = observations_to_image(result, {})
draw_ob = use_ob[:]
from PIL import Image
ob = Image.fromarray(draw_ob)
width, height = ob.size
[…]
return result, arm_action, end_ep
B.2 Dataset Details
(a) Task Classification
(b) Task Cloud
Figure B2: Dataset Statistics
In terms of task classification, the long-horizon task is the most, with 84 (76 in the train set and 8 in the test set). The least task is the most difficult open-ended task, with 6 in the train set and 3 in the test set. The distribution of the five types of tasks in the train and test sets is also approximately the same. In the word cloud map, we can see that put, find, sofa, etc. are popular words in our task.
B.3 Dataset Demonstration
Here we provide a demonstration of EMMOE data, its viusal information is shown in Fig. B3, the collection method is as Section 2.2.
Figure B3: Task Demonstration. line: subtask sequence; row: front, left, back, right
Task: check if there are anything in the kitchen drawer, if its empty put a short can into it
Subtask1: [Go to, blue_sofa] (PixNav)
Analysis: According to my final task, I must go to kitchen first, but I cant find kitchen in my sight and I find a blue sofa, so I should go to the blue sofa first and find kitchen later
Feedback: None
Inventory: None
Subtask2: [Go to, living_room] (NoMaD)
Analysis: I successfully arrived sofa but I still cant find kitchen, as I see living room, I should go to living room then keep searching kitchen
Feedback: None
Inventory: None
Subtask3: [Go to, kitchen_drawer] (NoMaD)
Analysis: I successfully arrived the living room and find the kitchen drawer, so I should go to kitchen drawer next
Feedback: None
Inventory: None
Subtask4: [Open, kitchen_drawer] (octo)
Analysis: I successfully arrived the kitchen drawer but its closed. To check items in the drawer, I should open it first
Feedback: None
Inventory: None
Subtask5: [Go to, kitchen_counter] (NoMaD)
Analysis: I successfully opened the drawer and find nothing in it, according to my task, I should put a short can into it in this case, as I find one on the kitchen counter, I should go there next
Feedback: None
Inventory: None
Subtask6: [Pick, short_can] (RT-1-X)
Analysis: I successfully arrived the kitchen counter and find the can, so I should pick it next
Feedback: None
Inventory: short_can
Subtask7: [Go to, kitchen_drawer] (NoMaD)
Analysis: I successfully picked the short can and I should come back to kitchen drawer to put it in
Feedback: None
Inventory: short_can
Subtask8: [Put, short_can, kitchen_drawer] (RT-1-X)
Analysis: I successfully arrived the kitchen drawer and I should put the can into it next
Feedback: None
Inventory: None
Subtask9: [Close, kitchen_drawer] (octo)
Analysis: I successfully put the can into the drawer, and its better to close the drawer next
Feedback: None
Inventory: None
Subtask10: [End]
Analysis: According to the historical execution and final task, I have finally finished the task and its time to end
Feedback: None
Inventory: None
We also provide all designed tasks here, the task design principles focus on reflecting human’s real-life with a variety of common demands and task descriptions.
(1) fetch a frozen meat can and put it on the kitchen counter
(2) clean up the brown table and place all items in the fridge
(3) find a cold apple and put it on the kitchen_counter
(4) find an bowl and put it on the sofa
(5) find an master_chef_can on the wood_table and put it into the drawer
(6) go to the floor 2
(7) prepare neccessary ingredients to make a fruit salad and put them on the yellow_table
(8) keep the number of red_boxes on the yellow_table to 5
(9) search a blue can for me
(10) fetch one crack box and one sugar box and put them on the beige table
(11) find two cracker boxes in the room and put them on the kitchen counter
(12) check if there are apples in the fridge and put one into it if not
(13) pick all fruit on the brown table and put them on the sofa
(14) put the bowl into the kitchen cabinet
(15) find a bleach cleanser and a sponge then place them on the brown table
(16) fetch two apples from the kitchen counter and put them into the fridge
(17) clean the wood table and put all items except mug to the sofa
(18) I want to eat at the brown table and prepare a fish can for me
(19) fetch two cracker_boxes from the kitchen sink and refrigerate them
(20) check and close all kitchen facilities
(21) prepare two bowls on the brown table
(22) fetch two meat_cans from the kitchen and put them on the beige table
(23) find a mug and put it on the tvstand
(24) go to kitchen then put the red box into the drawer and put the red can into the fridge
(25) find an apple and place it on the tv_stand
(26) clean the tvstand and put all items to the sofa
(27) clean up the tv_stand and put all items in the kitchen drawer
(28) put the sponge and bleach cleanser on the sofa into the kitchen drawer
(29) freeze a sugar_box
(30) put the blue can on the kitchen_counter to the fridge
(31) find two potted_meat_cans and put them on the sofa
(32) clean up the blue table and put all items to the white cabinet
(33) find an apple and put it on the sofa
(34) take a bowl and a meat can from the kitchen and put them on the brown table
(35) clean up the kitchen sink and put fruit to the fridge other items to the kitchen_counter
(36) replenish the number of blue cans on the table to 3
(37) find two bowls in the room and put them in the kitchen sink
(38) put all cracker_boxes on the tvstand to the sofa
(39) take a yellow box and put it into the fridge.
(40) put the apple on the blue table to the sofa
(41) fetch 3 different kinds of fruit and put them on the beige table
(42) I want to eat at the brown table and prepare some fruit for me
(43) put the frozen sponge into the kitchen drawer
(44) put all bowls on the sofa to the kitchen sink
(45) get a can in the fridge and put it on the table
(46) prepare a washed apple then put it on the yellow table
(47) clean up the tvstand
(48) clean up the chair
(49) put everything in the kitchen sink onto the kitchen_counter
(50) wash the bowl on the kitchen_counter
(51) fetch two sugar boxes in the fridge and put them on the brown table, if there arent enough sugar boxes in the fridge, find them elsewhere in the room
(52) Prepare a soup_can and a red_bowl on the kitchen_counter
(53) put all the fruit on the kitchen_counter into the sink
(54) put the bowl on the wood_table and the apple on the kitchen_counter to the kitchen sink
(55) refrigerate all master_chef_cans on the tvstand
(56) clean up the blue sofa
(57) find a gelation_box and put it in the drawer
(58) put the cracker box in the kitchen sink to the sofa
(59) check if there is food on the sofa then put them in the fridge if so
(60) refrigerate all lemons in the kitchen drawer
(61) put all food on the sofa into the drawer
(62) take the bowl on the table to the kitchen
(63) clean up the tv_stand and place items on the kitchen_counter
(64) check if there are bananas in the fridge; if not, get one from the kitchen and put it in the fridge
(65) fetch a yellow box from the refrigerator and place it on the table, if there isnt one, get it from the kitchen
(66) clean the sofa and put all items on the table in front of it
(67) find an apple and place it in the drawer
(68) Put the red bowl on the blue table in the fridge.
(69) go to the second floor
(70) keep the number of red_boxes on the yellow_table to 3 and put extra red_boxes to the kitchen_counter
(71) clean up the beige table and put all items to kitchen
(72) put all fruit in the living_room to the fridge
(73) find an apple and place it in the fridge
(74) find a bowl and a mug then put them into the kitchen sink
(75) replenish the number of pears in the fridge to 3
(76) find an apple and put it on the brown table
(77) put all lemons and apples on the sofa to the tvstand
(78) put all bowls in the open drawer onto the kitchen_counter
(79) clean up the sofa and put all items into the drawer
(80) clean up the sofa and place all items on the nearby chair
(81) freeze the meat can on the blue desk
(82) check and close all appliances in the room
(83) get a cold apple and put it on the wood table
(84) check if there are anything in the kitchen drawer, if its empty put a short can into it
(85) turn off all appliances in the room then go the door and wait
(86) prepare some food and put it on the brown table
(87) check items in the fridge then increase the number of blue cans to 2
(88) find a box and put it on the tvstand
(89) clean the table in front of you and put all items into the sink
(90) find two bananas on the tvstand and put them to the kitchen sink
(91) find the bowl in the drawer and put it to the kitchen sink
(92) get a cold fruit and prepare to wash it
(93) clean the sofa
(94) put all items on the sofa to the tvstand
(95) put all items on the blue sofa to the white desk
(96) find the sponge and put it into the drawer
(97) find two kinds of fruit and put them on the tvstand
(98) find a banana and place it in a bowl
(99) put the bowl on the brown table into the kitchen sink and put the suger_box on the tvstand to the sofa
(100) put the green_can on the brown_table to the fridge
Appendix C Metric Calculation
C.1 Task Progress
In the task demonstrated in Appendix B, it’s easy to find that to complete the task, we have to open the drawer to see if there is anything, and then we have to finish a put operation (put short can in the drawer). In addition to these two, we can also add some operation like, go to the drawer, close the cook and other actions which do not influence the final success. So we get the keypath as shown below,
"[Open, kitchen_drawer]",
"[Put, short_can, kitchen_drawer]",
"[End]"
"[Open, kitchen_drawer]",
"[Put, short_can, kitchen_drawer]",
"[Close, drawer]",
"[End]"
"[Go to, drawer]",
"[Open, kitchen_drawer]",
"[Put, short_can, kitchen_drawer]",
"[End]"
"[Go to, drawer]",
"[Open, kitchen_drawer]",
"[Put, short_can, kitchen_drawer]",
"[Close, drawer]",
"[End]"
Here’s an example to show how to calculate TP,
(1) [Go to, kitchen](success)
(2) [Open, drawer](success)
(3) [Put, short_can, drawer](fail)
(4) [Go to, kitchen_counter](success)
(5) [Put, short_can, kitchen_counter](fail)
(6) [Go to, drawer](success)
(7) [Put, short_can, drawer](fail)
(8) [Go to, kitchen_counter](success)
(9) [Put, short_can, kitchen_counter](fail)
(10) [Go to, drawer](success)
(11) [Put, short_can, drawer](fail)
(12) [Go to, kitchen_counter](success)
(13) [Put, short_can, kitchen_counter](fail)
(14) [Go to, drawer](success)
(15) [Put, short_can, drawer](fail)
(16) [Go to, kitchen_counter](success)
(17) [Put, short_can, kitchen_counter](fail)
(18) [Go to, drawer](success)
(19) [Put, short_can, drawer](fail)
(20) [Go to, kitchen_counter](success)
This is the result of one run, and we can see that the TP of this run is as calculated in Section 2.3, $max$ ($\frac{1}{3}$, $\frac{1}{4}$, $\frac{1}{2}$, $\frac{2}{5}$) = 0.5.
C.2 Success End Rate
In the above result, the number of steps reach 20, and there is no $End$ action to terminate the task. Here’s a example to show the success end.
(1) [Go to, kitchen_counter](success)
(2) [Go to, yellow_box](success)
(3) [Pick, yellow_box](success)
(4) [Go to, fridge](success)
(5) [Put, yellow_box, fridge](fail)
(6) [Open, fridge](fail)
(7) [Go to, kitchen_counter](success)
(8) [Put, yellow_box, kitchen_counter](success)
(9) [Go to, fridge](success)
(10) [Open, fridge](success)
(11) [Go to, kitchen_counter](success)
(12) [Pick, yellow_box](success)
(13) [Go to, fridge](success)
(14) [Put, yellow_box, fridge](success)
(15) [Close, fridge](success)
(16) [End]
This is the result of one run for the task take a yellow box and put it into the fridge, and we can judge by its keypath that it complete the task successfully. It has $End$ action, so the $End$ is a success end which can be treated as one of the numerators when calculating SER in Section 2.3. In fact, as said in Section 2.3, successful task trajectory must have one end, but there maybe other unsuccessful task trajectories have ends, that’s why we calculating SER.
C.3 Success Re-plan Rate
First of all, the next action our agent takes after the previous action failed is called replan. Use the above subsection result as an example, and it’s a successful task trajectory. In the step 5, the agent try to put the yellow box in the fridge but failed, and then, it try to open the fridge which can be treated as a success replan even though it failed again. Since the action open fridge is a meaningful action which can lead to the final success. It’s one of the numerators when calculating SRR in Section 2.3. Also, in the first subsection for TP, the example is an unsuccessful task trajectory, so actions like put short can drawer are not success replan.
Appendix D High Level Planning
In this section, we will should how the high-level planner described in Section 3.2 works step by step. A running demonstration of our HomieBot is shown in Fig. D4. To provide more intuitive understanding, we extract core sections from the original code and adapt them into a more general and easy-to-understand format to illustrate the process flow, this processing method is also applied to all subsequent code demonstrations. First, we provide the system information used in HomieBot, and all subsequent references to system information are consistent with what is provided here.
Figure D4: An illustration of running pipeline of HomieBot. After receiving images and feed- back, LMM generates analysis, specific subtask and model choice for low level executor to per- form.
You are a powerful housework assistant, I will give you following information for you to make a decision toward the final task.
(1) Observation images: Four first-person perspective images of the current environment, in the order of front, left, back, and right.
(2) Task: Your final goal.
(3) Inventory: Your current assets, remember that you are a one-hand agent, which means you cant open or pick when your Inventory is not None, and you cant put if your Inventory is None, this is very important.
(4) Historical Execution: Subtasks that were already fulfilled in the history, and the execution status of each subtask(success or fail). You need to make decisions based on historical actions, current circumstances and your final task.
(5) Feedback: Feedback will provide error information of the last execution, it will be None if the last execution ends successfully.
You should output with following formats:
Analysis: Make a detailed summary of your current situation based on given information, analyse and decide what to do next and output the reason of your decision.
Subtask: [action, target], choose your action from the action list [Go to, Pick, Put, Open, Close, End], and the target can be a place or a object from your observation. If you choose Put as your action, output in format [Put, object, place] which means put the object to the place. If the final task is done and no more action is needed, just output [End].
Model: Choose one most suitable model in the model list [NoMaD, PixNav, octo, RT-1-X]. NoMaD can go to a spot like living room, PixNav focuses on object navigation and can go to a object, octo can handle with open and close, RT-1-X is good at picking and putting.
You need to focus on the consistency with previous subtasks. You should pay attention to current Inventory and avoid conflicts.
Remember you can only go to the place and interact with the objects you observe in your sight.
Remember the logic between outputs, it is recommended to open the receptacle before you pick something because you cant open while holding, and its recommended to arrive the object place before you interact with it.
Remember you just need to output the next subtask to be fulfilled and dont output a whole plan, this is very important.
Remember you should output strictly with the response template.
Now, I will send the message so that you can make planning accordingly.
Next, we define some classes to make the overall process more readable and smooth. Here we only list most relevant and important parts in the process.
import os
import json
import re
class Conversations:
def __init__(self, max_round=20):
self.system = SYSTEM_INFO
self.history = []
self.round = 0
self.window = 3
self.max_round = max_round
def get_history_prompt(self):
history_prompt = ""
if self.round < self.window:
history_prompt = "".join(self.history)
else:
history_prompt = "".join(self.history[-3:])
return history_prompt
def reset(self):
self.history = []
self.round = 0
def save(self, save_path):
with open(os.path.join(save_path, "conversation.json"), "w") as file:
json.dump(self.history, file, indent=4)
class HomieBot:
def __init__(self):
self.conv = Conversations()
self.inventory = []
self.comm = Communicator()
def get_inventory(self):
if len(self.inventory) == 0:
return "None"
else:
return " ".join(self.inventory)
def generate_instruction(self, task, feedback, historical_execution):
if historical_execution == "":
[…]
else:
[…]
return instruction
def update_inventory(self, subtask, feedback):
subtask = subtask.lower()
if "None" in feedback:
if "pick" in subtask:
obj = subtask.split.split(’,’)[1].strip()
self.inventory.append(obj)
if "put" in subtask:
self.inventory.pop()
else:
if "put" in subtask and "the object is missing" in feedback:
self.inventory.pop()
def end(self):
self.comm.close_connection()
the most important function generate_instruction works as described in Section 3.2, which contains task, inventory, history and feedback.
Afterward, we provide the process for HomieBot to execute the task in a single trajectory.
homie = HomieBot()
task = "input your task"
save_path = "save_path"
feedback = ""
historical_execution = ""
while homie.conv.round < homie.conv.max_round:
homie.conv.round += 1
instruction = homie.generate_instruction(task, feedback, historical_execution)
images = homie.comm.receive_env_images()
output = model_inference(instruction, images)
[…]
pattern = r’.*Analysis: *(.+?) *Subtask: *\[(.*?)\].*Model: *(.*?)\$’
match = re.search(pattern, output, re.DOTALL)
analysis = match.group(1).strip()
subtask = match.group(2).strip()
model_choice = match.group(3).strip()
homie.comm.send_subtask(subtask, model_choice, homie.get_inventory())
feedback, signal = homie.comm.receive_feedback()
homie.update_inventory(subtask, feedback)
historical_execution += f"({homie.conv.round}) {subtask}({signal}) "
if "end" in subtask.lower():
break
homie.conv.save(save_path)
homie.end()
the realization of function model_inference varies from different models, but it’s quite easy to deploy different models into HomieBot as we can see in the code.
Appendix E Low Level Execution
E.1 Pipeline
def error_detection(action, target, inventory, env):
# Format Error Detection
if action not in action_list:
return ’fail’, f’{action} is not in the action list! You should only choose actions in the list.’
mapping_dict = load_name_mapping()
if target in mapping_dict:
target = mapping_dict[target]
else:
return ’fail’, f’{target} does not exist! Please choose another object’
# Logical Error Detection
if inventory != ’None’ and action in [’pick’, ’open’, ’close’]:
return ’fail’, f’Unable to {action}, the hand is full’
if inventory == ’None’ and action == ’put’:
return ’fail’, f’Unable to {action}, the hand is empty’
if action == ’put’ and "closed" in check_status(target):
return ’fail’, f’Unable to put, the {target} is closed, you should open it first’
if action in [’open’,’close’] and "non-interactive" in check_status(target):
return ’fail’, f’Can not {action} {target}! Please choose another object’
# Distance Error Detection
if action != "go to":
distance = calculate_distance(env, target)
if distance > 2:
return ’fail’, f’Unable to {action}, the target is far away’
if distance < 0.1:
return ’fail’, f’Unable to {action}, the target is too close’
return ’success’, ’None’
max_count = 20
comm = Communicator()
save_path = "save_path"
count_steps = 1
env = init_env()
while count_steps <= max_count:
images = get_env_images(save_path, env, count_steps)
comm.send_env_images(images)
action, target, inventory = comm.receive_subtask()
if "end" in action.lower():
comm.send_feedback("None", "success")
break
# Error Detection Before Execution
signal, feedback = error_detection(action, target, inventory, env)
if signal == "fail":
comm.send_feedback(feedback, signal)
break
for retry in range(3):
reset_arm(env)
# Error Detection During and After Execution
signal, feedback, env = execution(action, target, inventory, env)
if signal == ’success’:
break
elif action == ’put’ and env[’grasped_obj’] is None:
feedback = f’Unable to {action}, and the object is missing’
break
elif retry == 2:
feedback = f’Unable to {action}, the subtask is too difficult to perform’
if signal == ’success’:
feedback = "None"
count_steps += 1
comm.send_feedback(feedback, signal)
E.2 Skills
The skill we choose and their functions are shown in Table E1.
Table E1: The list of skills we used with descriptions and examples
Skill Description Example
Pick object Pick an object up pick sugar box
Put object to place Put an object into a place put lemon on brown table
Open container Open the container open the fridge
Close container Close the container close the kitchen drawer
Go to place navigate to a place navigate TV stand
Go to object navigate to where an object is navigate bowl
End End the execution End
E.3 Models
Table E2: Descriptions of Low Level Models used in HomieBot.
Model Input Capability Task
RT-1-X[67] RGB & Instructions Manipulation Picking & Placing
Octo[31] RGB & Instructions Manipulation Opening & Closing
NoMaD[32] RGB & Goal-Image Image-Navigation Navigate to Spot & Large Object
PixNav[33] RGB & Goal-Name Pixel-Navigation Navigate to Object
M3
[29] can flexible interact with target objects from various locations based on the integration of manipulative skills and mobility, while navigational skills are designed to accommodate multiple endpoints, ultimately leading to successful operations. Specifically, M3 implements these concepts by emphasizing mobile manipulation skills over fixed skills and training navigational skills using area targets rather than point targets.
RT-1-X
[30]) architecture utilizes image and text instructions as inputs, and generates discrete end-effector actions as outputs. Specifically, RT-1-X is a transformer-based model that guides robotic arms to complete various manipulation tasks. RT-1-X is an extension of the RT-1 ( [67]) model, which is designed for robot control and trained on a large-scale robot dataset.
Octo
[31]) is an open-source, general-purpose policy for robotic manipulation based on transformers. It supports flexible task and observation definition and can be quickly integrated into new observation and action spaces.
NoMaD
[32]) trains a single diffusion strategy for goal-oriented navigation and goal-independent exploration, the first one is to reach user-specified goals after localization and the second one is to search new environments. The method is instantiated using a transformer-based large-scale policy trained on data from various ground robots.
PixNav
[33]) is a pixel-guided navigational skill. It designs an LLM-based planner that utilizes common sense between objects and rooms to select the optimal waypoints, which are then executed by a pixel navigation strategy to achieve long-line-of-sight navigation. In this pipeline, we use its ability of finding the optimal waypoint and pixel navigation to navigate to some specific small object such as lemon and sugar box.
E.4 Error Classification
Logical error
If the hand already has an object (inventory is not empty) but still attempts to perform a pick/open/close operation, the execution will fail, and the message the hand is full will be returned; if the hand has no object (inventory is empty) but still attempts to perform a place operation, the execution will fail, and the message the hand is empty will be returned; if the item is not a container but still attempts to perform a open/close operation, the execution will fail, and the message please choose another object will be returned. In the execution with environment state information, if the container is closed and a place operation is still attempted, the execution will fail, and the message the container is closed, you should open it first will be returned.
Distance error
In the execution with environment state information, if the agent is too close to the target, causing the arm to be unable to extend properly but still attempts to perform a pick/place/open/close operation, the execution will fail, and the message the target is too close will be returned; if the agent is too far from the target, causing it to be unable to reach the target object but still attempts to perform a pick/place/open/close operation, the execution will fail, and the message the target is far away will be returned.
Format Error
For high level planning, it may output an object which is not in the scene, that is, in low level execution, we can’t find an object with a name matching the input in the scene, the message please choose another object will be returned; also, high level planning may output in a wrong operation which can not be performed, the message You should only choose actions in the list will be returned.
Execution Error
Due to the limited capabilities of low-level models, sometimes the failure is not caused by HLP. Therefore, each action can be executed up to three times. If it fails after three times, it will return a message the subtask is too difficult to perform; also, when performing a put operation, if the agent put the wrong place, it will return a message the object is missing to remind the agent to re-plan and re-pick.
Appendix F Data Augmentation
F.1 SFT Augmentation
To expand the original dataset size, we first use GPT-4o [8] to regenerate text descriptions. Here is the regeneration code clip, we just show how to regenerate task descriptions, but the regeneration of subtask analysis uses the same template.
client = OpenAI(api_key=’’)
completion = client.chat.completions.create(
model="gpt-4o",
messages=[
{"role": "system", "content": "Rewrite the following text with the same meaning but in a different description while do not change object’s name: "},
{"role": "user", "content": task}
Next we show how to convert a single EMMOE data into fix-format conversation data. After processing, each individual subtask will be combined with all previously subtasks to form a SFT data.
import os
with open(task_path) as file:
content = file.read()
[…]
task = content[0]
historical = ""
sft_data = []
for i, subtask_info in enumerate(content[1:]):
subtask_data = {}
[…]
if subtask_info[0] == ’’:
continue
subtask_id, decision = subtask_info[0].split(’: )
subtask_id = subtask_id.lower()
analysis = subtask_info[1]
if "End" not in decision:
action, model_choice = decision.strip(’)’).split( (’)
else:
action = "[End]"
model_choice = "None"
image_paths = [
os.path.join(save_dir, f"{subtask_id}_front.png"),
os.path.join(save_dir, f"{subtask_id}_left.png"),
os.path.join(save_dir, f"{subtask_id}_back.png"),
os.path.join(save_dir, f"{subtask_id}_right.png")
for path in image_paths:
if not os.path.exists(path):
raise FileNotFoundError(f"File does NOT exist: {path}")
if i == 0:
[…]
else:
[…]
feedback = subtask_info[2]
inventory = subtask_info[3]
if "None" in feedback:
[…]
else:
[…]
conv = []
conv.append({"from": "human", "value": instruction})
conv.append({"from": "gpt", "value": answer})
task_id = ’_’.join(task_path.split(’.’)[0].split(’/’)[-2:])
subtask_data["id"] = task_id + ’_’ + subtask_id
subtask_data["image"] = image_paths
subtask_data["conversations"] = conv
sft_data.append(subtask_data)
We also provide some data samples for more intuitive understanding.
"id": "21_info_subtask13",
"image": [
"./HomieBot/data/train/21/subtask13_front.png",
"./HomieBot/data/train/21/subtask13_left.png",
"./HomieBot/data/train/21/subtask13_back.png",
"./HomieBot/data/train/21/subtask13_right.png"
"conversations": [
"from": "human",
[…]
"from": "gpt",
[…]
"id": "21_info_subtask14",
"image": [
"./HomieBot/data/train/21/subtask14_front.png",
"./HomieBot/data/train/21/subtask14_left.png",
"./HomieBot/data/train/21/subtask14_back.png",
"./HomieBot/data/train/21/subtask14_right.png"
"conversations": [
"from": "human",
[…]
"from": "gpt",
[…]
F.2 DPO Augmentation
The last step for SFT Augmentation and the first step of DPO Augmentation occur simultaneously, failed subtasks are considered junk data for SFT but can be used to construct DPO data, so we run $data\_sift$ function first. After getting initial DPO data, we will use SFT data, $model\_change$, $order\_change$ and $action\_change$ to construct final DPO dataset.
def data_sift(subtask_list):
sft_data = []
dpo_data = []
flag = 1
for i in range(1, len(subtask_list)):
if "Feedback: None" in subtask_list[i]["conversations"][0]["value"]:
sft_data.append(subtask_list[i-1])
if flag == 0:
dpo_data.append({
"prompt": subtask_list[i-2]["conversations"][0]["value"],
[…]
flag = 1
else:
flag = 0
sft_data.append(subtask_list[-1])
return sft_data, dpo_data
def dpo_augment(sft_data, dpo_data):
for i in range(len(sft_data)):
prompt = sft_data[i]["conversations"][0]["value"]
[…]
if "End" in sft_data[i]["conversations"][1]["value"]:
continue
def model_change(chosen):
if "NoMaD" in chosen:
return chosen.replace("NoMaD", "PixNav")
elif "PixNav" in chosen:
return chosen.replace("PixNav", "NoMaD")
elif "octo" in chosen:
return chosen.replace("octo", "RT-1-X")
else:
return chosen.replace("RT-1-X", "octo")
def order_change(i, sft_data):
[…]
def action_change(chosen):
if "Pick" in chosen:
return chosen.replace("Pick", "Fetch")
elif "Put" in chosen:
return chosen.replace("Put", "Place")
elif "Go to" in chosen:
return chosen.replace("Go to", "Move")
elif "Open" in chosen:
return chosen.replace("Open", "Pull")
elif "Close" in chosen:
return chosen.replace("Close", "Push")
reject1 = model_change(chosen)
reject2 = order_change(i, sft_data)
reject3 = action_change(chosen)
dpo_data.append({"prompt": prompt, "chosen": chosen, "rejected": reject1})
dpo_data.append({"prompt": prompt, "chosen": chosen, "rejected": reject2})
dpo_data.append({"prompt": prompt, "chosen": chosen, "rejected": reject3})
return dpo_data
Notably, action $End$ is special among all available actions and it will only appear as $rejected$ in DPO data. In the first augmentation stage and $order\_change$, since the relationship between $chosen$ and $rejected$ is $O_{i}$ and $O_{i+1}$ (see definitions in Section 4.1) and there are no other subtasks after $End$, which means other actions might appear in either $chosen$ or $rejected$ while $End$ can only be the $rejected$. But this effect of suppressing the $End$ output is exactly what we want. Even executing a few extra steps after completing the task is better than terminating early without finishing the task. That is to say, We hope the model could consider more and do not output $End$ so easily. Experimental results in Table 2 and Table 3 confirm the effectiveness of this method as we can see an improvement in $SER$ metric, another positive phenomenon in results is that the length of the successful paths hasn’t increased significantly as we observe in $PLWSR$ and $TP$.
Finally, we provide some DPO data examples.
[…]
Appendix G Training Details
G.1 Training Parameters
We use Video-LLaVA-7B [110] as our base model, we also use the training scripts they provide and partial parameters for $sft$ are as follows.
--lora_enable True
--lora_r 128
--lora_alpha 256
--mm_projector_lr 2e-5
--bits 4
--mm_projector_type mlp2x_gelu
--mm_vision_select_layer -2
--mm_use_im_start_end False
--mm_use_im_patch_token False
--image_aspect_ratio pad
--group_by_modality_length True
--bf16 True
--num_train_epochs 1
--per_device_train_batch_size 16
--per_device_eval_batch_size 4
--gradient_accumulation_steps 1
--evaluation_strategy "no"
--save_strategy "steps"
--save_steps 50000
--save_total_limit 1
--learning_rate 5e-4
--weight_decay 0.
--warmup_ratio 0.03
--lr_scheduler_type "cosine"
--logging_steps 1
--tf32 True
--model_max_length 2048
--tokenizer_model_max_length 3072
--gradient_checkpointing True
--dataloader_num_workers 4
--lazy_preprocess True
--report_to tensorboard
We use finetuned model as our base and reference model, and use open-source $trl$ package and parameters for $dpo$ are as follows.
bnb_config = BitsAndBytesConfig(
load_in_4bit=True,
bnb_4bit_compute_dtype=torch.float16,
bnb_4bit_use_double_quant=True,
bnb_4bit_quant_type=’nf4
training_args = DPOConfig(
per_device_train_batch_size=16,
per_device_eval_batch_size=4,
gradient_accumulation_steps=1,
gradient_checkpointing=True,
max_grad_norm=0.3,
num_train_epochs=1,
save_steps=1000,
learning_rate=5e-6,
bf16=True,
save_total_limit=1,
logging_steps=10,
output_dir=output_dir,
optim="paged_adamw_32bit",
lr_scheduler_type="cosine",
warmup_ratio=0.03,
remove_unused_columns=False
peft_config = LoraConfig(
r=8,
lora_alpha=8,
target_modules=find_all_linear_names(model),
lora_dropout=0.05,
bias="none",
task_type="CAUSAL_LM",
dpo_trainer = DPOTrainer(
model,
model_ref,
args=training_args,
beta=0.1,
train_dataset=train_dataset,
eval_dataset=eval_dataset,
tokenizer=tokenizer,
max_prompt_length=2048,
max_length=2048,
Appendix H Experimental Details
H.1 Baseline Setup
To make it more convenient for different models to deploy into our system without training, we slightly lower output format requirements, here shows the adapatations.
import re
pattern = r’.*Analysis: *(.+?) *Subtask: *\[(.*?)\].*Model: *(.*?)\$’
match = re.search(pattern, output, re.DOTALL)
if match == None:
pattern = r’.*Analysis: *(.+?) *Subtask: *(.*?) *Model: *(.*?)\$’
match = re.search(pattern, output, re.DOTALL)
Despite lowering the output format standards, the output from 7B-sized models still fails to meet our least requirements. They either do not output single-step subtasks or the subtask format is far from requirements. This issue is difficult to resolve by merely adjusting prompts. Therefore, we leverage the in-context learning abilities of these models by providing an output template example before each inference. Here, we provide the inference template for Qwen2-VL [39] MiniCPM-V 2.6 [40] respectively.
Qwen2VL
messages = [
{"role": "system", "content": homie.conv.system},
{"role": "user",
[…]
{"role": "assistant",
"content": "I will surely follow the given format, now you can send prompt to me."
{"role": "user",
"content": [
{"type": "image", "image": images[0]},
{"type": "image", "image": images[1]},
{"type": "image", "image": images[2]},
{"type": "image", "image": images[3]},
{"type": "text", "text": instruction}]
prompt = processor.apply_chat_template(
messages, tokenize=False, add_generation_prompt=True
image_inputs, video_inputs = process_vision_info(messages)
inputs = processor(
text=[prompt],
images=image_inputs,
videos=video_inputs,
padding=True,
return_tensors="pt"
)to("cuda")
generated_ids = model.generate(**inputs, max_new_tokens=512)
enerated_ids_trimmed = [
out_ids[len(in_ids) :] for in_ids, out_ids in zip(inputs.input_ids, generated_ids)
outputs = processor.batch_decode(
generated_ids_trimmed, skip_special_tokens=True, clean_up_tokenization_spaces=False
MiniCPM-V 2.6
image_loads = [Image.open(image).convert(’RGB’) for image in images]
messages = [
{"role": "user",
[…]
{"role": "assistant",
"content": "I will surely follow the given format, now you can send prompt to me.",
{"role": "user",
"content": [image_loads[0], image_loads[1], image_loads[2], image_loads[3], instruction]
output = model.chat(
image=None,
system_prompt=homie.conv.system,
tokenizer=tokenizer
H.2 Results
Here we provide more detailed results of experiments in Section 4.5. Table H3 and Table H4 show the statistics results in percentages while Table H5 and Table H6 show original counts. Table H7 show the original counts and success rate range of each action.
Table H3: Successful Trajectories Error Statistics All definitions are same as in Section 4.5. Additionally, we add statistics of four primary types.
Models L1 L2 L3 L4 L D1 D2 D F1 F2 F E1 E2 E All
GPT-4o[8] 3.97 0.79 0.79 0 5.56 44.44 0 44.44 1.59 17.46 19.05 15.87 15.08 30.95 30.29
Gemini-1.5-Pro[37] 3.85 3.85 0 7.69 15.38 48.08 0 48.08 0 17.31 17.31 15.38 3.85 19.23 21.80
Qwen2-VL-7B[39] 0 0 0 0 0 100 0 100 0 0 0 0 0 0 20
MiniCPM-V 2.6[40] 0 0 0 0 0 100 0 100 0 0 0 0 0 0 6.67
o1[38] 0 0 1.18 10.06 11.24 21.89 0 21.89 0 48.52 48.52 18.34 0 18.34 20.56
HomieBot-7B (SFT) 10.53 9.77 12.78 1.50 34.59 36.09 0 36.09 0 3.01 3.00 24.06 2.26 26.32 14.41
HomieBot-7B (SFT+DPO) 10.17 15.25 9.32 3.39 38.14 33.05 0 33.05 0 3.39 3.39 25.42 0 25.42 12.87
Table H4: Failed Trajectories Error Statistics
Models L1 L2 L3 L4 L D1 D2 D F1 F2 F E1 E2 E All
GPT-4o[8] 6.87 0.12 0.69 3.65 11.34 8.41 0.06 8.47 0.57 64.88 65.45 13.99 0.75 14.74 73.61
Gemini-1.5-Pro[37] 7.48 1.52 2.41 6.45 17.86 9.41 0 9.41 0 47.86 47.86 22.76 2.10 24.86 68.38
Qwen2-VL-7B[39] 2.17 9.49 0.99 3.56 16.21 7.71 0 7.71 4.74 54.35 59.09 16.40 0.59 17.00 27.74
MiniCPM-V 2.6[40] 8.58 0.80 0.92 1.72 12.01 7.78 0 7.78 3.49 65.39 68.88 10.87 0.46 11.33 31.08
o1[38] 1.16 0.07 1.96 4.86 8.05 8.64 0 8.64 0 56.75 56.75 26.27 0.29 26.59 51.77
HomieBot-7B (SFT) 11.31 23.85 9.86 4.20 49.24 11.77 0 11.77 0.61 11.47 12.08 24.54 2.37 26.91 35.70
HomieBot-7B (SFT+DPO) 11.46 23.90 11.13 2.62 49.10 9.25 0 9.25 0.25 17.27 17.51 22.67 1.47 24.14 35.88
Table H5: Original Successful Trajectories Statistics All data are integers.
Models L1 L2 L3 L4 L D1 D2 D F1 F2 F E1 E2 E All
GPT-4o[8] 5 1 1 0 7/126 56 0 56/126 2 22 24/126 20 19 39/126 126/416
Gemini-1.5-Pro[37] 4 4 0 8 16/104 50 0 50/104 0 18 18/104 16 4 20/104 104/477
Qwen2-VL-7B[39] 0 0 0 0 0/9 9 0 9/9 0 0 0/9 0 0 0/9 9/45
MiniCPM-V 2.6[40] 0 0 0 0 0/1 1 0 0/1 0 0 0/1 0 0 0/1 1/15
o1[38] 0 0 2 17 19/169 37 0 37/169 0 82 82/169 31 0 31/169 169/822
HomieBot-7B (SFT) 14 13 17 2 46/133 48 0 48/133 0 4 4/133 32 3 35/133 133/923
HomieBot-7B (SFT+DPO) 12 18 11 4 45/118 39 0 39/118 0 4 4/118 30 0 30/118 118/917
Table H6: Original Failed Trajectories Statistics
Models L1 L2 L3 L4 L D1 D2 D F1 F2 F E1 E2 E All
GPT-4o[8] 228 4 23 121 376/3317 279 2 281/3317 19 2152 2171/3317 464 25 489/3317 3317/4506
Gemini-1.5-Pro[37] 217 44 70 187 518/2900 273 0 273/2900 0 1388 1388/2900 660 61 721/2900 2900/4241
Qwen2-VL-7B[39] 33 144 15 54 246/1518 117 0 117/1518 72 825 897/1518 249 9 258/1518 1518/5472
MiniCPM-V 2.6[40] 150 14 16 30 210/1748 136 0 136/1748 61 1143 1204/1748 190 8 198/1748 1748/5624
HomieBot-7B (SFT) 148 312 129 55 644/1308 154 0 154/1308 8 150 158/1308 321 31 352/1308 1308/3664
HomieBot-7B (SFT+DPO) 140 292 136 32 600/1222 113 0 113/1222 3 211 214/1222 277 18 295/1222 1222/3406
Table H7: $Count$ represents the number of each action error with a total count behind. $SR_{range}$ is the percentage to indicate the range of success rates of each action, with the average value shown in parentheses.
Metrics Go to Pick Place Open Close
Count 938/2437 1213/2437 178/2437 81/2437 27/2437
$SR_{range}$ (45.32)31.19$\sim$82.89 (22.45)18.34$\sim$28.95 (40.97)29.06$\sim$69.40 (43.13)30.13$\sim$75.84 (36.45)26.72$\sim$57.37
H.3 Task Performance
We also evaluate SR for each type of task defined in Section 2.2. As shown in Table H8, typical tasks are relatively easy due to straightforward processes and fewer overall steps. The most challenging are open-ended tasks, which usually have a very long total step count, with flexible processes and results, demanding powerful capabilities from both HLP and LLE models.
Table H8: Average success rate for each type of task. The highest value of each model is highlighted in bold.
Model Short- Horizon Long- Horizon Open- Ended Logical Human- Style
HomieBot (SFT) 43.75 24.60 18.52 34.01 25.24
HomieBot (SFT+DPO) 41.67 28.11 15.38 35.86 27.88
Appendix I Case Study
We show case studies of the inference from HomieBot in various situations as follows.
Case 1: Successful trajectory
Here we show a successful trajectory of our HomieBot (DPO version). To facilitate understanding, we convert the dialogue data into the original EMMOE data format. As shown, even if errors occur, timely adjustments can be made through feedback, ensuring the correctness of the execution process.
Task: put the green_can on the brown_table to the fridge
#Round1
Analysis: According to my final task, I need to go to the green_can first.
Subtask: [Go to, green_can]
Inventory: None
Feedback: None
#Round2
Analysis: I successfully arrived the green_can and I should pick it next
Subtask: [Pick, green_can]
Inventory: green_can
Feedback: None
#Round3
Analysis: I successfully picked the green_can, so according to my final task, I should go to the fridge to put it down.
Subtask: [Go to, fridge]
Inventory: green_can
Feedback: None
#Round4
Analysis: I successfully arrived the fridge and I should put the green_can into it next
Subtask: [Put, green_can, fridge]
Inventory: green_can
Feedback: Unable to put, the fridge is closed, you should open it first
#Round5
Analysis: I failed to put the green_can into the fridge, since the fridge is closed, so I should open the fridge first
Subtask: [Open, fridge]
Inventory: green_can
Feedback: Unable to open, the hand is full
#Round6
Analysis: I failed to open the fridge because my hand is full, so I should put the green_can on the brown_table first
Subtask: [Go to, brown_table]
Inventory: green_can
Feedback: None
#Round7
Analysis: I successfully arrive the brown_table and I should put the green_can on it
Subtask: [Put, green_can, brown_table]
Inventory: None
Feedback: None
#Round8
Analysis: I successfully put the green_can on the brown_table, so I should go to the fridge next
Subtask: [Go to, fridge]
Inventory: None
Feedback: None
#Round9
Analysis: I successfully arrived the fridge and I should open it next
Subtask: [Open, fridge]
Inventory: None
Feedback: None
#Round10
Analysis: I successfully open the fridge and find the green_can, so I should go to the green_can.
Subtask: [Go to, green_can]
Inventory: None
Feedback: None
#Round11
Analysis: I successfully arrive the green_can and find it, so I should pick it.
Subtask: [Pick, green_can]
Inventory: green_can
Feedback: None
#Round12
Analysis: I successfully picked the green_can, so according to my final task, I should go to the fridge to put it down.
Subtask: [Go to, fridge]
Inventory: green_can
Feedback: None
#Round13
Analysis: I successfully arrived the fridge and I should put the green_can into it next
Subtask: [Put, green_can, fridge]
Inventory: None
Feedback: None
#Round14
Analysis: I successfully put the green_can into the fridge, so I should close the fridge next
Subtask: [Close, fridge]
Inventory: None
[…]
#Round15
Analysis: According to the historical execution and final task, I have finally finished the task and its time to end
Subtask: [End]
Inventory: None
Feedback: None
Case 2: Terrible Grounding
Here we show a terrible grounding problem issues during the inference of GPT4-o [8]. Due to the lengthy path, we only highlight the most critical subtask outputs and their execution results to emphasize errors (following cases are handled similarly). We can see that the ungrounded output directly prevents the process to be continued. Even after informing the model that the object doesn’t exist, the issue remains unresolved. The model continues to output incorrect objects or makes mistakes again after a few steps.
Task: clean up the beige table and put all items to kitchen
(1) Go to, beige table(success)
(2) Pick, item(fail)
(3) Pick, item(fail)
(4) Pick, item(fail)
(5) Pick, item(fail)
(6) Pick, item(fail)
(7) Pick, item(fail)
(8) Pick, item on the table(fail)
(9) Pick, item on the table(fail)
(10) Pick, item on the table(fail)
(11) Pick, item on the table(fail)
(12) Pick, item(fail)
(13) Pick, item on the table(fail)
(14) Pick, item on the beige table(fail)
(15) Pick, stack of papers(fail)
(16) Pick, box(fail)
(17) Go to, closer to the table(fail)
(18) Pick, item on the table(fail)
(19) Pick, visible item on the beige table(fail)
Case 3: Limited LLE
Here we show a problem issues from the limited ability of low-level models. As we can see in step7 and step8, though high-level planner makes correct plans, execution still fails due to the limited ability of low-level models, this problem occurs repeatedly during whole process and finally leads to the failue the trajectory.
Task: take a yellow box and put it into the fridge.
(1) Go to, living room(success)
(2) Pick, yellow box(success)
(3) Go to, kitchen(success)
(4) Open, fridge(fail)
(5) Put, yellow box, counter(fail)
(6) Put, yellow box, table(fail)
(7) Go to, fridge(success)
(8) Open, fridge(fail)
(9) Put, yellow box, counter(fail)
(10) Open, fridge(fail)
(11) Put, yellow box, counter(fail)
(12) Open, fridge(fail)
(13) Put, yellow box, counter(fail)
(14) Open, fridge(fail)
(15) Put, yellow box, counter(fail)
(16) Open, fridge(fail)
(17) Put, yellow box, counter(fail)
(18) Open, fridge(fail)
(19) Put, yellow box, counter(fail)
Case 4: Meaningless Outputs
Here we show a case of Qwen2-VL [39] generating a large amount of meaningless outputs during the inference. As we can see, even though all subtasks are successful, the agent keep circling in place without making progress. These meaningless outputs quickly consume the remaining execution steps, ultimately causing the task to fail.
Task: find two bowls in the room and put them in the kitchen sink
(1) Go to, kitchen(success)
(2) Go to, kitchen(success)
(3) Go to, kitchen(success)
(4) Go to, kitchen(success)
(5) Go to, kitchen(success)
(6) Go to, kitchen(success)
(7) Go to, living room(success)
(8) Go to, living room(success)
(9) Go to, kitchen(success)
(10) Go to, kitchen(success)
(11) Go to, living room(success)
(12) Go to, living room(success)
(13) Go to, living room(success)
(14) Go to, living room(success)
(15) Go to, living room(success)
(16) Go to, living room(success)
(17) Go to, living room(success)
(18) Go to, living room(success)
(19) Go to, living room(success)
Case 5: Solvable D1 Error
Here we show how $D1$ error is solvable during the inference of Gemini-1.5-Pro [37]. As we can see in step3, after a $D1$ error happens, a $Go\ to$ action can effectively solve it and facilitate the success of the whole trajectory.
Task: fetch one crack box and one sugar box and put them on the beige table
(1) [Go to, beige_table](success)
(2) [Pick, cracker_box](fail)
(3) [Go to, cracker_box](success)
(4) [Pick, cracker_box](success)
(5) [Go to, beige_table](success)
(6) [Put, cracker_box, beige_table](success)
(7) [Go to, brown_table](success)
(8) [Put, sugar_box, brown_table](fail)
(9) [Go to, beige_table](success)
(10) [Put, sugar_box, beige_table](fail)
(11) [Go to, sugar_box](success)
(12) [Pick, sugar_box](success)
(13) [Go to, beige_table](success)
(14) [Put, sugar_box, beige_table](success)
(15) [End](success)
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: In Section 2, we propose EMMOE, the first unified benchmark designed to evaluate both high-level planners and low-level policies. In Section 2.2, we present the collection and features of EMMOE-100. In Section 2.3, we propose three novel metrics to complement existing evaluation methods. Next, we introduce our HomieBot and illustrate how its two main components HLP and LLE function in Section 3. In Section 4, we conduct experiments and demonstrate how to construct LMM-trainable SFT and DPO datasets and evaluate different levels of models. Finally, we conduct an in-depth analysis based on the detailed error information in Section 4.5.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: In Section 5, we discuss following limitations and future works: Limited actions and available space in Habitat restrict the scope of task design. Besides, standardized output will sacrifice certain information precision. The growing number of model inferences will also lead to additional time costs. Moreover, we choose to conduct evaluations solely in simulation, where all researchers are required to make assessments under the same conditions, thus ensuring optimal fairness and consistency. In the future, we’ll collect more tasks, design a more efficient system, and explore real-world evaluations.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We propose three new metrics in our work, and we have provide detailed definitions in Section 2.3, visible calculation processes in Appendix C, and complete evaluation codes in supplementary materials and project website.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: To ensure the reproducibility of our work, we provide data examples and visualizations in Appendix B and F, detailed training parameters and partial codes in Appendix H, entire codes in supplementary materials and project website. We also upload our models and datasets to the Huggingface platform.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We have uploaded our codes on Github, our models and dataset on Huggingface, demonstrations on Youtube. we provide a project website at the end of the abstract, where all links can be found there.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, c, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: In Section 4.2, we provide some essential hyperparameter during model training, and full training settings in Appendix G. In Section 4.3, we introduce our metrics, baseline models, training and testing split selection.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification:
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: In Section 4.2 and Section 4.3, we discuss the type of compute workers GPU of model training and experiments.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have carefully checked the codes we provide, to avoid potential harms and and better serve for the community.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: In Section 3, we also discuss the potential benefits of our system and setting. Though we evaluate on simulation, our methods have the potentials to facilitate to the real-world deployments and running. In Section 5, we discuss the limitations of our work.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Our data is collected in an open-sourced simulator, and our models are also trained based on the open-source models. We have carefully follow the instructions of these simulators and models.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We have included citations of all works we used and mentioned and clearly mentioned the license and terms of use, which can be checked in the main paper and Appendix A.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We have provided detailed and complete instructions and Readme file of our dataset, model, and codes in our supplementary materials and project website.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: Our work doesn’t involve human participation, and doesn’t talk about the human-involved case either.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Our work does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: Our work uses LLM for training and evaluations, as well a part of our proposed system. We have clearly pointed out what LLMs we use and their explicit functions in our work. We have also reported their performance and detailed analysis during experiments.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.